OpenStudy (anonymous):

Solve for b2 in A = h(b1+b2), if A = 16, h = 4, and b1 = 3.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

in the formula, divide both sides by h; then subtract b1 from both sides... this solves for b2... now just plug in those givens to get a value for b2....:)

OpenStudy (anonymous):

16=4(3+b2) 16=12+4b2 16-12=4b2 4=4b2 4/4=b2 b2=1

OpenStudy (anonymous):

plug b2=1 into the equation to test.